Grade 12th PassAnalytical Geometry

If x=√ [(2+√3)/(2-√3)], then the value of x2(x-4)2 is?

Charitha Gorrela
9 Years agoGrade 12th Pass

1 Answer

Vikas TU
9 Years ago
.X=√ [(2+√3)/(2-√3)], 
At that point x2(x-4)2= 
Justifying x=√ [(2+√3)/(2-√3)] with √ (2+√3) 
We get x=√7+4√3 
Presently x2=7+4√3 
(x-4)2=(√7+4√3-4)2 
=23+4√3-8√7+4√3 
Presently x2(x-4)2=161+28√3-56√7+4√3+92√3+48+24√21+12√3 
=209+120√3-56√7+4√3+24√21+12√3 .